# FedHK-MVFC: Federated Heat Kernel Multi-View Clustering

### v1

@@ 3.7 HK-MVFC Algorithm
Algorithm HK-MVFC provides a detailed implementation of the HK-MVFC in Eq. 6 approach. The algorithm operates by alternately updating the membership matrix, cluster centers, and view weights until convergence.
@@
10:    Compute $\delta_{ij}^{h^{(t)}}$ using Eq. 1 or 2
@@ 3.8 Computational Complexity of HK-MVFC
The computation of heat-kernel coefficients $\delta_{ij}^{h}$ using either Eq. 1 or 2 represents the initial computational overhead. For each view, this calculation requires $O(nd_{h})$ operations across all $n$ data points. When extended across all $s$ views, the total complexity for heat-kernel coefficient computation becomes $O(nsd_{h})$ per iteration.
@@ 4 Federated HK-MVFC
To implement HK-MVFC within a federated setting, we must reformulate the objective function for distributed optimization across multiple clients while preserving data privacy and ensuring convergence guarantees. Let $\mathcal{M}=\{1,2,\ldots,M\}$ denote the set of $M$ participating clients, where each client $\ell\in\mathcal{M}$ maintains a local multi-view dataset $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ with potentially heterogeneous view compositions and sample distributions. Here, $s(\ell)$ denotes the number of views for client $\ell$, and may vary across clients, reflecting the possibility that each client possesses a different number of data views.
@@ 4.1 Problem Formulation
Consider a federated setting with $M$ clients, where each client $\ell$ holds $s(\ell)$ views of local data. For client $\ell$, the $h$-th view is denoted as $X_{[\ell]}^{h}=\{x_{[\ell]i}^{h}\}_{i=1}^{n(\ell)}$, where $x_{[\ell]i}^{h}\in\mathbb{R}^{d_{[\ell]}^{h}}$ is the $i$-th sample in view $h$. Thus, the complete multi-view dataset for client $\ell$ is $X_{[\ell]}=\{X_{[\ell]}^{1},X_{[\ell]}^{2},\ldots,X_{[\ell]}^{s(\ell)}\}$, with $X_{[\ell]}^{h}\in\mathbb{R}^{n(\ell)\times d_{[\ell]}^{h}}$. Each client possesses the same feature set for each view, but the number of samples per view may vary across clients. Definitions include:
  • $n(\ell)$ denotes the quantity of data samples possessed by client $\ell$
  • $s(\ell)$ refers to the number of views client $\ell$ holds
  • $d_{\left[\ell\right]}^{h}$ signifies the dimensionality of the $h$-th view for client $\ell$
@@ 4.2 Federated KED
We assume that while participating clients may contribute varying numbers of samples, each data view shares the same set of features. Thus, the actual or expected cluster formations of multi-view datasets managed by the clients are not required to be identical. This implies that clients can handle multi-view datasets with differing quantities of clusters. By considering this, we can extend $\text{KED}_{2}\left({x_{ij}^{h},a_{kj}^{h}}\right)$ in Eq. 5 to a federated context, referred to as federated kernel Euclidean distance (FKED). The FKED is formulated as follows
@@ 4.3 Federated Heat-Kernel Coefficients
We focus on developing federated heat-kernel coefficients (FedH-KC), which are computed locally by each client to transform feature distances into privacy-preserving exponential kernel measures tailored to the federated setting. For the $j$-th feature in view $h$, we define FedH-KC $\delta_{\left[\ell\right]ij}^{h}$ using two distinct estimators.
@@
$\text{FedH-KC}=\delta_{\left[\ell\right]ij}^{h}=\frac{x_{\left[\ell\right]ij}^{h}-\min_{1\leq i\leq n(\ell)}\left(x_{\left[\ell\right]ij}^{h}\right)}{\max_{1\leq i\leq n(\ell)}\left(x_{\left[\ell\right]ij}^{h}\right)-\min_{1\leq i\leq n(\ell)}\left(x_{\left[\ell\right]ij}^{h}\right)+\epsilon}\qquad\forall j,h,\ell$ (22)
@@
where $\epsilon>0$ prevents division by zero when features have identical values. This normalizes features within each view for consistent distance calculations across clients with different data distributions. and
$\text{FedH-KC}=\delta_{\left[\ell\right]ij}^{h}=\left|{x_{\left[\ell\right]ij}^{h}-{{\bar{x}}_{\left[\ell\right]j}}^{h}}\right|\qquad\forall j,h,\ell$ (23)
where $\bar{x}_{[\ell]j}^{h}=\frac{1}{n(\ell)}\sum_{i=1}^{n(\ell)}x_{[\ell]ij}^{h}$ represents the mean value of the $j$-th feature in view $h$ for client $\ell$, harmonizing with the centralized case notation.
@@ 4.4 The Objective Function
In a federated setting with $M$ clients, we connect FedHK-MVFC and FKED before introducing the main objective function. The FedHK-MVFC algorithm extends the centralized HK-MVFC by distributing computation across clients while retaining heat kernel-enhanced distance properties. Each client $\ell$ holds local multi-view data $X_{[\ell]}$ and computes local heat-kernel coefficients $\delta_{[\ell]ij}^{h}$, converting Euclidean distances into geometry-aware similarities. Through the FKED, local client computations align with the federated objective. For client $\ell$, the local goal reflects the centralized HK-MVFC structure tailored to client-specific data, with local federated kernel distance $d_{\exp([\ell],ik,j)}^{h}=\text{FKED}\left(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h}\right)$ maintaining geometric properties within distributed processing. This federated system combines local results for global clustering coherence without losing data locality. On this basis, the FedHK-MVFC objective for client $\ell$ is defined as:
@@ 4.5 Optimization Framework
For the federated objective function $J_{FedHK-MVFC}^{\ell}$ defined in Eq. 24, the necessary conditions for optimality yield the following update rules for client $\ell$:
@@
where the updates are performed iteratively until convergence, subject to the normalization constraints in Eqs. 25 and 26.
@@
To derive the update equation for membership coefficients $\mu_{[\ell]ik}^{*}$ for client $\ell$, we begin by computing the partial derivative of the Lagrangian in Eq. 27 with respect to $\mu_{[\ell]ik}^{*}$.
@@
To determine the Lagrange multiplier $\lambda_{1i}$, we apply the normalization constraint $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$:
@@
Substituting this back into Eq. 35:
@@
This establishes the membership matrix update rule in Eq. 28.
@@
For cluster center updates $a_{[\ell]kj}^{h}$, we differentiate the objective function in Eq. 24 with respect to $a_{[\ell]kj}^{h}$:
@@
Let $\phi_{[\ell]ik}^{h}=\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}$. Then:
@@
Substituting this into Eq. 40 and setting equal to zero for optimality:
@@
Simplifying and rearranging:
@@
Let $w_{[\ell]ijk}^{h}=(\mu_{[\ell]ik}^{*})^{m}v_{[\ell]h}^{\alpha}\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$. Then:
@@
This yields the cluster center update rule in Eq. 29.
@@
For view weight updates $v_{[\ell]h}$, we differentiate the Lagrangian with respect to $v_{[\ell]h}$:
@@
This establishes the view weight update rule in Eq. 30.
@@
    Stationarity: $\nabla J_{FedHK-MVFC}^{\ell}+\sum_{i=1}^{n(\ell)}\lambda_{1i}\nabla g_{i}+\lambda_{2}\nabla h=0$, where $g_{i}(\mu)=\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}-1$ and $h(v)=\sum_{h=1}^{s(\ell)}v_{[\ell]h}-1$.
@@
    Primal feasibility: $g_{i}(\mu)=0$ and $h(v)=0$ for all $i$.
@@
    Dual feasibility: All Lagrange multipliers are finite and well-defined.

### v2

@@ 3.7 HK-MVFC Algorithm
Algorithm 1 provides a detailed implementation of the HK-MVFC in Eq. 13 approach. The algorithm operates by alternately updating the membership matrix, cluster centers, and view weights until convergence.
@@
10:    Compute $\delta_{ij}^{h^{(t)}}$ using Eq. 8 or 9
@@
3.7.1 HK-MVFC Flowchart
Figure 1 presents the complete workflow of the centralized Heat Kernel-Enhanced Multi-View Fuzzy Clustering algorithm, detailing the initialization, iterative optimization, and convergence checking phases.
[…]
Figure 1: Flowchart of the HK-MVFC Algorithm (Algorithm 1). The algorithm iteratively updates membership matrix, cluster centers, and view weights using heat kernel-enhanced distances until convergence.
@@ 3.8 Computational Complexity of HK-MVFC
The computation of heat-kernel coefficients $\delta_{ij}^{h}$ using either Eq. 8 or 9 represents the initial computational overhead. For each view, this calculation requires $O(nd_{h})$ operations across all $n$ data points. When extended across all $s$ views, the total complexity for heat-kernel coefficient computation becomes $O(nsd_{h})$ per iteration.
@@ 4 Federated HK-MVFC
To implement HK-MVFC within a federated setting, we must reformulate the objective function for distributed optimization across multiple clients while preserving data privacy and ensuring convergence guaranties. Let $\mathcal{M}=\{1,2,\ldots,M\}$ denote the set of $M$ participating clients, where each client $\ell\in\mathcal{M}$ maintains a local multi-view dataset $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ with potentially heterogeneous view compositions and sample distributions. Here, $s(\ell)$ denotes the number of views for client $\ell$, and may vary across clients, reflecting the possibility that each client possesses a different number of data views.
@@ 4.1 Problem Formulation
Consider a federated learning environment comprising $M$ participating clients, where each client $\ell\in\{1,2,\ldots,M\}$ maintains its own local multi-view dataset. The fundamental characteristic of this federated setting is that data remains distributed across clients, with no centralized data repository, thereby preserving data privacy and institutional autonomy.
4.1.1 Multi-View Data Structure at Each Client
For each client $\ell$, the local dataset consists of $s(\ell)$ distinct views, where each view captures a different modality or perspective of the underlying phenomenon. The $h$-th view at client $\ell$ is mathematically represented as:
$X_{[\ell]}^{h}=\{x_{[\ell]i}^{h}\}_{i=1}^{n(\ell)}\subset\mathbb{R}^{d_{[\ell]}^{h}}$ (28)
where $x_{[\ell]i}^{h}\in\mathbb{R}^{d_{[\ell]}^{h}}$ denotes the $i$-th data sample in the $h$-th view, characterized by $d_{[\ell]}^{h}$ features. This representation can be organized as a matrix:
$X_{[\ell]}^{h}\in\mathbb{R}^{n(\ell)\times d_{[\ell]}^{h}}$ (29)
where rows correspond to individual samples and columns represent features within that specific view.
The complete multi-view dataset for client $\ell$ encompasses all available views:
$X_{[\ell]}=\{X_{[\ell]}^{1},X_{[\ell]}^{2},\ldots,X_{[\ell]}^{s(\ell)}\}$ (30)
This formulation allows for heterogeneity in the number of views across clients, reflecting realistic scenarios where different institutions may collect different types of complementary data.
4.1.2 Notation and Parameter Definitions
To ensure mathematical precision and clarity, we establish the following notation conventions:
  • Sample Count: $n(\ell)\in\mathbb{Z}^{+}$ represents the total number of data samples (e.g., patients, observations) available at client $\ell$. This parameter may vary significantly across clients, reflecting differences in institutional size, data collection capacity, or temporal duration of data acquisition.
  • View Count: $s(\ell)\in\mathbb{Z}^{+}$ denotes the number of distinct views or data modalities maintained by client $\ell$. For instance, in a medical scenario, one hospital might possess $s(\ell)=2$ views (ECG and MRI data), while another might have $s(\ell)=3$ views (ECG, MRI, and genetic data).
  • Feature Dimensionality: $d_{[\ell]}^{h}\in\mathbb{Z}^{+}$ specifies the dimensionality (number of features) of the $h$-th view at client $\ell$. For example, an ECG view might have $d_{[\ell]}^{1}=12$ (corresponding to 12 leads), while an MRI view might have $d_{[\ell]}^{2}=1024$ (representing voxel features).
  • Cluster Count: $c(\ell)\in\mathbb{Z}^{+}$ indicates the number of target clusters for client $\ell$. While clients may share a common clustering objective (e.g., $c(\ell)=c$ for all $\ell$), our framework accommodates scenarios where different clients seek to identify different numbers of local clusters.
4.1.3 Key Properties and Assumptions
The federated multi-view clustering framework operates under several important properties and assumptions:
  1. 1.
    Feature Space Consistency: While clients may possess different numbers of samples ($n(\ell)$ varies), we assume that corresponding views across clients share the same feature space dimensionality: $d_{[\ell]}^{h}=d_{[\ell^{\prime}]}^{h}$ for all clients $\ell,\ell^{\prime}\in\{1,\ldots,M\}$ that possess view $h$. This ensures meaningful aggregation of model parameters across clients.
  2. 2.
    Data Locality: Each client’s data $X_{[\ell]}$ remains strictly local and is never transmitted to other clients or the central server. Only model parameters (cluster centers, membership matrices, view weights) are communicated, ensuring privacy preservation.
  3. 3.
    Statistical Heterogeneity: The data distributions may differ significantly across clients: $P_{\ell}(X_{[\ell]})\neq P_{\ell^{\prime}}(X_{[\ell^{\prime}]})$ for $\ell\neq\ell^{\prime}$. This non-IID (non-independent and identically distributed) property reflects realistic federated scenarios where different institutions serve different populations or geographic regions.
  4. 4.
    View Complementarity: The multiple views at each client provide complementary rather than redundant information. Mathematically, for views $h$ and $h^{\prime}$ at client $\ell$:
    $\text{MI}(X_{[\ell]}^{h},X_{[\ell]}^{h^{\prime}})<H(X_{[\ell]}^{h})\quad\text{and}\quad\text{MI}(X_{[\ell]}^{h},X_{[\ell]}^{h^{\prime}})<H(X_{[\ell]}^{h^{\prime}})$ (31)
    where $\text{MI}(\cdot,\cdot)$ denotes mutual information and $H(\cdot)$ represents entropy, ensuring that each view contributes unique information to the clustering task.
4.1.4 Federated Learning Objective
The overarching goal in this federated setting is to collaboratively learn a global clustering model that effectively integrates information across all clients and views, while respecting the following constraints:
  • Privacy Preservation: Raw data $X_{[\ell]}$ never leaves client $\ell$
  • Communication Efficiency: Minimize the volume of data transmitted between clients and server
  • Clustering Quality: Achieve clustering performance comparable to centralized approaches where all data would be pooled
  • Geometric Awareness: Capture the intrinsic manifold structure of multi-view data through heat kernel-enhanced distance measures
This problem formulation establishes the mathematical foundation for developing the FedHK-MVFC algorithm, which addresses these objectives through a principled integration of federated learning, multi-view clustering, and heat kernel methods.
@@ 4.2 Federated KED
We assume that, while participating clients may contribute varying numbers of samples, each data view shares the same set of features. Thus, the actual or expected cluster formations of multi-view datasets managed by the clients are not required to be identical. This implies that clients can handle multi-view datasets with differing quantities of clusters. By considering this, we can extend $\text{KED}_{2}\left({x_{ij}^{h},a_{kj}^{h}}\right)$ in Eq. 12 to a federated context, referred to as the federated kernel Euclidean distance (FKED). The FKED is formulated as follows:
@@ 4.3 Federated Heat-Kernel Coefficients
We focus on developing federated heat-kernel coefficients (FedH-KC), which are computed locally by each client to transform feature distances into privacy-preserving exponential kernel measures tailored to the federated setting. Before presenting the mathematical formulations, we establish the theoretical connection between quantum field theory (QFT) and heat-kernel coefficients that motivates our approach.
4.3.1 Quantum Field Theory Foundation for Heat-Kernel Coefficients
The heat-kernel coefficients employed in our framework originate from the heat equation in quantum field theory, which describes the diffusion of information across a manifold. In QFT, the heat kernel $K_{t}(\mathbf{x},\mathbf{y})$ represents the fundamental solution to the heat equation on a Riemannian manifold $\mathcal{M}$:
$\left(\frac{\partial}{\partial t}+\Delta_{\mathcal{M}}\right)K_{t}(\mathbf{x},\mathbf{y})=0,\quad K_{0}(\mathbf{x},\mathbf{y})=\delta(\mathbf{x}-\mathbf{y})$ (33)
where $\Delta_{\mathcal{M}}$ denotes the Laplace-Beltrami operator on the manifold, and $\delta(\cdot)$ is the Dirac delta function. The heat kernel admits the spectral decomposition:
$K_{t}(\mathbf{x},\mathbf{y})=\sum_{i=0}^{\infty}e^{-\lambda_{i}t}\phi_{i}(\mathbf{x})\phi_{i}(\mathbf{y})$ (34)
where $\{\lambda_{i},\phi_{i}\}$ are the eigenvalue-eigenfunction pairs of the Laplacian operator, encoding the intrinsic geometry of the data manifold.
4.3.2 From QFT to Practical Heat-Kernel Coefficients
The connection between QFT heat kernels and our practical coefficients emerges through the asymptotic expansion of the heat kernel for small diffusion time $t$:
$K_{t}(\mathbf{x},\mathbf{y})\sim(4\pi t)^{-d/2}e^{-\frac{d^{2}(\mathbf{x},\mathbf{y})}{4t}}\sum_{n=0}^{\infty}t^{n}a_{n}(\mathbf{x},\mathbf{y})$ (35)
where $d(\mathbf{x},\mathbf{y})$ represents the geodesic distance on the manifold, $d$ is the manifold dimension, and $\{a_{n}\}$ are the heat-kernel coefficients capturing local geometric invariants such as curvature and torsion.
For practical clustering applications on discrete data, we approximate the continuous heat kernel using the exponential distance kernel:
$K(\mathbf{x}_{i},\mathbf{x}_{j})\approx\exp\left(-\sum_{h=1}^{s}\delta_{ij}^{h}\|\mathbf{x}_{i}^{h}-\mathbf{x}_{j}^{h}\|^{2}\right)$ (36)
where the coefficients $\delta_{ij}^{h}$ serve as adaptive scaling factors that encode local geometric properties analogous to the heat-kernel coefficients $a_{n}$ in the continuous theory. These coefficients effectively modulate the diffusion rate based on local data density and feature variance, thereby capturing the intrinsic manifold structure.
4.3.3 Geometric Interpretation and Information Diffusion
The heat-kernel framework provides a natural interpretation of clustering as information diffusion on the data manifold. Points within the same cluster exhibit rapid information exchange (high heat kernel values), while points in different clusters experience slow diffusion across cluster boundaries. The federated heat-kernel coefficients $\delta_{[\ell]ij}^{h}$ quantify this diffusion rate locally at each client, enabling:
  1. 1.
    Geometry-Aware Similarity: By normalizing distances according to local feature scales, the coefficients ensure that similarity measures respect the intrinsic geometric structure rather than arbitrary coordinate choices.
  2. 2.
    Manifold Structure Preservation: The exponential decay in Eq. (36) mirrors the heat kernel’s behavior on curved manifolds, where geodesic distances differ from Euclidean distances.
  3. 3.
    Multi-Scale Analysis: Different choices of normalization (min-max vs. mean-variance) correspond to different time scales in the heat diffusion process, enabling multi-scale geometric analysis.
This theoretical foundation justifies our use of heat-kernel coefficients in federated multi-view clustering, where each client’s local geometric structure must be captured while maintaining global clustering coherence.
4.3.4 Federated Heat-Kernel Coefficient Formulations
Building upon the QFT foundation, we now present two practical estimators for federated heat-kernel coefficients. For the $j$-th feature in view $h$, we define FedH-KC $\delta_{[\ell]ij}^{h}$ using:
Type 1 - Min-Max Normalization (Uniform Scaling):
@@ 4.3.4 Federated Heat-Kernel Coefficient Formulations
$\text{FedH-KC}_{1}=\delta_{[\ell]ij}^{h}=\frac{x_{[\ell]ij}^{h}-\min_{1\leq i\leq n(\ell)}\left(x_{[\ell]ij}^{h}\right)}{\max_{1\leq i\leq n(\ell)}\left(x_{[\ell]ij}^{h}\right)-\min_{1\leq i\leq n(\ell)}\left(x_{[\ell]ij}^{h}\right)+\epsilon}\qquad\forall j,h,\ell$ (37)
@@
where $\epsilon>0$ prevents division by zero when features have identical values. This normalizes features within $[0,1]$ for each view, corresponding to uniform heat diffusion across the feature space.
Type 2 - Mean-Variance Normalization (Adaptive Scaling):
$\text{FedH-KC}_{2}=\delta_{[\ell]ij}^{h}=\left|x_{[\ell]ij}^{h}-\bar{x}_{[\ell]j}^{h}\right|\qquad\forall j,h,\ell$ (38)
where $\bar{x}_{[\ell]j}^{h}=\frac{1}{n(\ell)}\sum_{i=1}^{n(\ell)}x_{[\ell]ij}^{h}$ represents the mean value of the $j$-th feature in view $h$ for client $\ell$. This formulation adapts to local data distributions, corresponding to diffusion rates proportional to local variance.
Both formulations preserve the essential heat-kernel property of geometry-aware distance transformation while maintaining computational efficiency and privacy in the federated setting. The choice between Type 1 and Type 2 depends on the specific geometric characteristics of the data and the desired sensitivity to local versus global structure.
@@ 4.4 The Objective Function
In a federated setting with $M$ clients, we connect FedHK-MVFC and FKED before introducing the main objective function. The FedHK-MVFC algorithm extends the centralized HK-MVFC by distributing computation across clients while retaining the heat kernel-enhanced distance properties. Each client $\ell$ holds local multi-view data $X_{[\ell]}$ and computes local heat-kernel coefficients $\delta_{[\ell]ij}^{h}$, converting Euclidean distances into geometry-aware similarities through the lens of heat diffusion on the data manifold.
Through the FKED, local client computations align with the federated objective. For client $\ell$, the local goal reflects the centralized HK-MVFC structure tailored to client-specific data, with the local federated kernel distance $d_{\exp([\ell],ik,j)}^{h}=\text{FKED}\left(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h}\right)$ maintaining geometric properties within distributed processing. This federated system combines local results for global clustering coherence without losing data locality. On this basis, the FedHK-MVFC objective for client $\ell$ is defined as:
@@ 4.5 Optimization Framework
For the federated objective function $J_{FedHK-MVFC}^{\ell}$ defined in Eq. 39, the necessary conditions for optimality yield the following update rules for client $\ell$. These update rules extend the centralized HK-MVFC framework to the federated setting by incorporating client-specific data distributions, heat-kernel coefficients, and privacy-preserving mechanisms while maintaining mathematical rigor and convergence guarantees.
@@
Interpretation: Eq. 43 computes the fuzzy membership degree $\mu_{[\ell]ik}^{*}$ of data point $i$ to cluster $k$ at client $\ell$. The numerator aggregates the weighted federated kernel distances across all $s(\ell)$ views available at client $\ell$, where $v_{[\ell]h}^{\alpha}$ represents the adaptive importance of view $h$, and the exponent $\frac{1}{m-1}$ controls the fuzziness of the partition. The denominator normalizes these memberships across all $c(\ell)$ clusters, ensuring that $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$ for each data point. This formulation enables the algorithm to automatically balance contributions from multiple heterogeneous views while adapting to local data characteristics at each client.
@@
Interpretation: Eq. 44 updates the $j$-th feature of cluster center $k$ in view $h$ for client $\ell$ as a weighted average of the corresponding data point features. The weight for each data point combines three critical components: (1) the fuzzified membership degree $(\mu_{[\ell]ik}^{*})^{m}$, which emphasizes points strongly belonging to cluster $k$; (2) the view importance $v_{[\ell]h}^{\alpha}$, which prioritizes more informative views; and (3) the exponential heat kernel term $\exp(-\delta_{[\ell]ij}^{h}\|x_{[\ell]i}^{h}-a_{[\ell]k}^{h}\|^{2})$, which provides geometry-aware weighting based on the local manifold structure captured by the federated heat-kernel coefficients $\delta_{[\ell]ij}^{h}$. This composite weighting mechanism ensures that cluster centers are positioned to reflect both the fuzzy partition structure and the intrinsic geometric properties of the local data at each client.
@@
Interpretation: Eq. 45 computes the adaptive weight $v_{[\ell]h}$ for view $h$ at client $\ell$ based on the clustering quality within that view. The numerator evaluates the total weighted distance of view $h$, where lower distances indicate better clustering performance and should result in higher view weights. The exponent $\frac{1}{\alpha-1}$ creates an inverse relationship between distance and weight, while the parameter $\alpha>1$ controls the sharpness of this relationship—larger values of $\alpha$ create more pronounced differences between high-quality and low-quality views. The denominator normalizes these weights across all $s(\ell)$ views to ensure $\sum_{h=1}^{s(\ell)}v_{[\ell]h}=1$. This adaptive weighting mechanism enables FedHK-MVFC to automatically identify and emphasize the most informative views at each client, accommodating heterogeneous data quality and view-specific noise characteristics across the federated network.
Remark 2.
The federated update rules in Eqs. 4345 preserve the mathematical structure of the centralized HK-MVFC framework while incorporating client-specific parameters $n(\ell)$, $s(\ell)$, and $c(\ell)$ that reflect data heterogeneity across the federation. The iterative alternating optimization—sequentially updating memberships, cluster centers, and view weights—converges to a local optimum satisfying the Karush-Kuhn-Tucker (KKT) conditions for the constrained federated optimization problem, subject to the normalization constraints in Eqs. 40 and 41. Crucially, all computations occur locally at each client using only their private data $X_{[\ell]}$, with only the resulting model parameters $(U_{[\ell]}^{*},A_{[\ell]},V_{[\ell]})$ being shared with the federated server for secure aggregation, thereby ensuring privacy preservation throughout the learning process.
@@
To derive the update equation for membership coefficients $\mu_{[\ell]ik}^{*}$ for client $\ell$, we begin by computing the partial derivative of the Lagrangian in Eq. 42 with respect to $\mu_{[\ell]ik}^{*}$. This derivative represents the marginal change in the objective function when adjusting the membership of data point $i$ to cluster $k$ at client $\ell$.
@@
In Eq. 47, the term $m(\mu_{[\ell]ik}^{*})^{m-1}$ arises from applying the power rule to the fuzzified membership term $(\mu_{[\ell]ik}^{*})^{m}$. The summation $\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})$ represents the weighted contribution of all views to the distance between data point $i$ and cluster center $k$, where $v_{[\ell]h}^{\alpha}$ adaptively weights each view based on its clustering quality.
@@
This stationarity condition ensures that the membership assignment minimizes the weighted distance objective while satisfying the normalization constraint.
@@
The exponent 1/(m1)1/(m-1) transforms the equation from its derivative form back to the original membership variable. The negative sign in the Lagrange multiplier term ensures that memberships decrease as weighted distances increase, which is the desired clustering behavior where closer points receive higher memberships.
To determine the Lagrange multiplier $\lambda_{1i}$, we apply the normalization constraint $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$, which ensures that the total membership of data point $i$ across all clusters sums to unity:
@@
The denominator in Eq. 52 serves as a normalization factor that ensures all memberships sum to one, effectively balancing the contributions from all clusters based on their respective distances to the data point.
Substituting Eq. 52 back into Eq. 50:
@@
This final form shows that each membership is inversely proportional to the weighted distance raised to the power 1/(m1)1/(m-1), with the proportionality constant chosen to satisfy the normalization constraint. This establishes the membership matrix update rule in Eq. 43.
@@
For cluster center updates $a_{[\ell]kj}^{h}$, we differentiate the objective function in Eq. 39 with respect to $a_{[\ell]kj}^{h}$. This derivative measures how the objective function changes when we adjust the $j$-th feature of cluster center $k$ in view $h$ for client $\ell$:
@@
where $v_{[\ell]h}^{\alpha}$ weights the contribution of view $h$, and $(\mu_{[\ell]ik}^{*})^{m}$ weights each data point’s contribution by its fuzzy membership to cluster $k$.
@@
Let $\phi_{[\ell]ik}^{h}=\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}$ represent the weighted squared Euclidean distance in the heat kernel exponent. The heat-kernel coefficients $\delta_{[\ell]ij^{\prime}}^{h}$ adapt to local data geometry, enabling the distance metric to capture intrinsic manifold structure. Then:
@@
The factor of 2 arises from differentiating the squared term $(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})^{2}$, while the exponential term weights contributions based on the heat kernel similarity between data points and cluster centers. The term $(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})$ indicates the direction of adjustment: positive when data points are larger than the center, negative otherwise.
Substituting Eq. 60 into Eq. 55 and setting equal to zero for optimality:
@@
Simplifying and rearranging by canceling constant factors:
@@
Let $w_{[\ell]ijk}^{h}=(\mu_{[\ell]ik}^{*})^{m}v_{[\ell]h}^{\alpha}\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$ denote the composite weight for data point $i$ in cluster $k$ for feature $j$ in view $h$. This weight combines fuzzy memberships $(\mu_{[\ell]ik}^{*})^{m}$, view importance $v_{[\ell]h}^{\alpha}$, and heat kernel similarities. Then:
@@
This final expression shows that cluster centers are weighted averages of data points, where the weights account for membership degrees, view importance, and geometric proximity via heat kernel similarities. This yields the cluster center update rule in Eq. 44.
@@
For view weight updates $v_{[\ell]h}$, we differentiate the Lagrangian with respect to $v_{[\ell]h}$. This derivative quantifies how the objective function changes when we adjust the importance assigned to view $h$ at client $\ell$:
@@
The term $\alpha v_{[\ell]h}^{\alpha-1}$ comes from differentiating $v_{[\ell]h}^{\alpha}$ with respect to $v_{[\ell]h}$ using the power rule, while the double summation $\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})$ represents the total weighted distance contribution of view $h$ across all data points and clusters.
@@
The exponent $1/(\alpha-1)$ transforms the derivative relationship back to the view weight variable. The negative exponent on the distance summation ensures that views with smaller total distances (better clustering quality) receive higher weights, implementing an adaptive mechanism that prioritizes more informative views.
@@
This adaptive weighting mechanism automatically assigns higher importance to more informative views, where informativeness is measured by the view’s ability to create compact, well-separated clusters. This establishes the view weight update rule in Eq. 45.
@@
    Stationarity: $\nabla J_{FedHK-MVFC}^{\ell}+\sum_{i=1}^{n(\ell)}\lambda_{1i}\nabla g_{i}+\lambda_{2}\nabla h=0$, where $g_{i}(\mu)=\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}-1$ and $h(v)=\sum_{h=1}^{s(\ell)}v_{[\ell]h}-1$. This condition ensures that the gradient of the Lagrangian vanishes at the optimal solution, accounting for both the objective function and the constraints.
@@
    Primal feasibility: $g_{i}(\mu)=0$ and $h(v)=0$ for all $i$. This guarantees that the solution satisfies all equality constraints, ensuring valid probability distributions for memberships and view weights.
@@
    Dual feasibility: All Lagrange multipliers are finite and well-defined. The multipliers $\lambda_{1i}$ and $\lambda_{2}$ exist and are uniquely determined by the normalization constraints, as shown in Eqs. 52 and the corresponding view weight derivation.